OpenStudy (anonymous):

what is the complete balanced equation for this reaction? CaCl2 + Na3PO4 yields NaCl + Ca3(PO4)2

OpenStudy (anonymous):

3CaCl2 + 2Na3PO4 = 6NaCl + Ca3(PO4)2

OpenStudy (anonymous):

3 CaCl2 + 2 Na3Po4 = Ca3(Po4)2 + 6 NaCl